### v2

Abstract
Class-agnostic counting (CAC) aims to estimate the number of objects in images without being restricted to predefined categories. However, while current exemplar-based CAC methods offer flexibility at inference time, they still rely heavily on labeled data for training, which limits scalability and generalization to many downstream use cases. In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that exploits a fully unsupervised feature extractor. Specifically, our approach employs self-supervised vision-only backbones to extract object-aware features, and it eliminates the need for annotated data throughout the entire proposed pipeline. At inference time, we extract latent object prototypes via ROI-Align from DINO features and use them as convolutional kernels to generate similarity maps. These are then transformed into density maps through a simple yet effective normalization scheme. We evaluate our approach on the FSC-147 benchmark, where we consistently outperform a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. Additionally, we achieve competitive results – and in some cases surpass – training-free methods that rely on supervised backbones, non-training-free unsupervised methods, as well as several fully supervised SOTA approaches. This demonstrates that label- and training-free CAC can be both scalable and effective. Website: https://lorebianchi98.github.io/CountingDINO/.
1 Introduction
Class-agnostic counting (CAC) removes the need for class-specific retraining by allowing users to define target categories at inference [DBLP:conf/cvpr/RanjanSNH21]. This recent paradigm addresses the inherent limitations of class-specific counting approaches, which rely on predefined object types, such as vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], cells [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16], or animals [DBLP:conf/eccv/ArtetaLZ16, DBLP:journals/cea/TianGCWLM19].
However, CAC models still rely heavily on annotated training data. First, most existing approaches extract image features using backbones pre-trained on large-scale labeled datasets such as ImageNet [DBLP:conf/cvpr/HeZRS16]. Second, they require supervised training on annotated datasets, including thousands of dot annotations and bounding boxes, limiting scalability. Dot annotations, placed on object centroids, are used to generate density maps that serve as training targets for a regression network learning to map image features to object density estimates [DBLP:conf/nips/LempitskyZ10]. Bounding boxes, on the other hand, are used to localize a small set of visual prototypes, known as exemplars, which represent the object category to be counted [DBLP:conf/cvpr/RanjanSNH21]. Particularly, the requirement for thousands of dot annotations poses a major obstacle to creating large-scale datasets. As a result, only a few datasets are available, and even these often exhibit notable limitations [10944113].
In this paper, we introduce CountingDINO, the first training-free exemplar-based CAC framework that eliminates the need for labeled data at any stage. Our image feature extraction backbone builds on DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], which is trained in a self-supervised fashion. At inference time, we extract exemplar features by applying ROI-Align [8372616] to user-provided bounding boxes using the same DINO-computed image feature maps. These exemplar features are then used as depthwise convolutional kernels over the image features, generating similarity maps that highlight regions matching the exemplars. We average these similarity maps across exemplars to produce a global, informative response, which is then converted into a density map through a simple yet effective normalization scheme. Final object counts are computed by integrating the density map. Furthermore, to address the limitations of the spatial resolution of DINO with small objects, we partition the image into non-overlapping quadrants, apply feature extraction independently, and aggregate the resulting maps. We validate our method on the gold-standard CAC benchmark FSC-147 [DBLP:conf/cvpr/RanjanSNH21], consistently outperforming a baseline based on an SOTA unsupervised object detector under the same label- and training-free setting. We also compare it to training-free methods relying on supervised backbones, non-training-free unsupervised methods, and fully supervised approaches, demonstrating that our method remains competitive despite not using supervision or training at any stage.
To summarize, we propose the following contributions:
  • We propose CountingDINO, the first training-free class-agnostic counting framework that uses self-supervised backbones to extract object-aware features, removing any need for annotated data or any form of training.
  • We assess our approach on the FSC-147 benchmark. CountingDINO outperforms a baseline under the same label- and training-free setting and achieves results that are comparable to – and in some cases surpass – recent training-free methods with supervised backbones, unsupervised approaches (that are not training-free), as well as fully supervised approaches.
  • We conduct ablation studies to assess the contribution of each core component in our model.
2 Related Works
2.1 Exemplar-based Class-agnostic Counting
Counting category-specific objects is a longstanding task in computer vision with broad real-world applications – e.g., counting of people [DBLP:conf/cvpr/LiuSF19, DBLP:journals/eswa/BenedettoCCFGA22], vehicles [DBLP:conf/iccv/ZhangWCM17, DBLP:conf/iscc/AmatoCFG19], insects [DBLP:journals/ecoi/CiampiZICBFAC23, DBLP:journals/cea/Bereciartua-Perez22], or biological structures [DBLP:journals/mia/CiampiCTMLSAPG22, DBLP:conf/eccv/XueRHB16]. Among existing counting approaches, density map regression – estimating counts via feature-to-density mapping – has proven effective in crowded scenes [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10], outperforming detection-based techniques [DBLP:journals/cea/TianGCWLM19, DBLP:conf/iscc/AmatoCFG19]. However, their reliance on category-specific annotations and training limits scalability and generalization.
Class-agnostic counting (CAC) has recently shifted the object counting paradigm toward open-world contexts, enabling models to handle arbitrary categories unseen during training. In this setting, users are no longer restricted to predefined categories and can instead specify novel object classes at inference time by providing visual exemplars – typically in the form of three bounding boxes enclosing visual prototypes within the same input image [DBLP:conf/cvpr/RanjanSNH21]. The seminal FamNet [DBLP:conf/cvpr/RanjanSNH21] combines multi-scale feature extraction with exemplar-image correlation via inner product operations for density map prediction. RCAC [DBLP:conf/eccv/GongZ0DS22] follows a similar architecture but introduces a feature augmentation module that synthesizes exemplars with varied colors, shapes, and scales to enhance diversity. BMNet [DBLP:conf/cvpr/Shi0FL022] highlights the limitations of inner product-based correlation and instead proposes a learnable bilinear similarity loss, inspired by metric learning, for better matching. LOCA [DBLP:conf/iccv/EukicLZK23] introduces an object prototype extraction module that iteratively refines exemplar features via cross-attention, while PseCo [DBLP:conf/cvpr/HuangD0ZS24] relies on the popular SAM for instance segmentation [DBLP:conf/iccv/KirillovMRMRGXW23]. CACViT [DBLP:conf/aaai/WangX0024] leverages a single pre-trained Vision Transformer, using its attention mechanism for both feature extraction and matching. DAVE [DBLP:conf/cvpr/PelhanLZK24] proposes a two-stage detect-and-verify framework: first identifying candidate boxes via density maps and then verifying them using exemplar-based clustering. In contrast to these supervised methods, UnCounTR [knobel2024uncountr] is trained without manual annotations by using self-collages generated from unsupervised data and DINO features. Instead, TFPOC [DBLP:conf/wacv/Shi0Z24] and OmniCount [DBLP:journals/corr/abs-2403-05435] propose training-free pipelines using SAM without point-level supervision — though SAM is label-trained, making the approach not fully label-free. A recent survey of CAC methods is available in [DBLP:journals/corr/abs-2501-19184].
2.2 Unsupervised Vision Backbones
Recent advances in unsupervised, especially self-supervised, learning have produced models with highly effective representations for classification and localization. The DINO family [caron2021emerging, oquab2023dinov2, darcet2024vision], based on vision transformers [dosovitskiy2021an] and trained via self-distillation, demonstrates that patch-level features learned without labels can capture rich semantic information, outperforming CLIP-based models, which rely on weak supervision from image-text pairs. Alongside DINO, masked autoencoders (MAE) [he2022masked] offer a self-supervised approach by training transformers to reconstruct masked patches, fostering global visual understanding.
In this work, we propose the first training-free exemplar-based CAC framework that leverages fully unsupervised vision backbones to extract features from both images and exemplars. Our pipeline requires no labeled data or supervision at any stage, neither for training the image feature extractor nor for generating the density maps.
3 Method
Given an image and a small set of exemplar bounding boxes from the same image – each containing a visual prototype of the object we want to count – we aim to estimate the total number of instances of that object class present in the image, without relying on human-labeled data at any stage. To this end, we leverage the object understanding capabilities of DINO [caron2021emerging, oquab2023dinov2, darcet2024vision], a self-supervised vision-only backbone, and proceeding as described in the following sections. We also report a detailed schema in Fig. 1.
3.1 DINO-based Feature Extraction
Let $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ be an input image, and let $\psi_{\text{v}}$ denote a DINO-based self-supervised visual backbone. Passing the image through $\psi_{\text{v}}$ yields a dense feature map: $\mathbf{V}=\psi_{\text{v}}(\mathbf{I})\in\mathbb{R}^{L\times V\times D},$ where $L=\frac{H}{P}$ and $V=\frac{W}{P}$ represent the spatial dimensions of the feature map, with $P$ being the patch size and $D$ the dimensionality of the feature embeddings. Each location in $\mathbf{V}$ encodes a feature vector corresponding to a specific image patch.
We also extract the representation from all the available exemplars. Each exemplar is specified by a bounding box in the original image space, defined by its top-left and bottom-right coordinates as $\mathbf{b}=(x_{1},y_{1},x_{2},y_{2})$, where $x_{1},y_{1},x_{2},y_{2}\in\mathbb{R}$. We simply perform ROI-Align on $\mathbf{V}$ using the bounding box $\mathbf{b}$, by carefully downsampling the coordinates of $\mathbf{b}$ to $\mathbf{\tilde{b}}=(\lfloor Lx_{1}\rfloor,\lfloor Vy_{1}\rfloor,\lceil Lx_{2}\rceil,\lceil Vy_{2}\rceil)$. The output of this operation is a feature $\mathbf{R}^{\mathbf{\tilde{b}}}$ that captures the visual appearance of the exemplar $\mathbf{\tilde{b}}$ in the embedding space.
To refine the exemplar representation, we apply a soft spatial prior in the form of an elliptical weighting mask centered within the bounding box. This mask attenuates peripheral regions and emphasizes the center under the assumption that exemplar objects are typically centered within their boxes. Formally, we define a mask $\mathbf{M}^{\mathbf{\tilde{b}}}\in\mathbb{R}^{w(\mathbf{\tilde{b}})\times h(\mathbf{\tilde{b}})}$, where each entry $m^{\mathbf{\tilde{b}}}_{ij}$ indicates the proportion of the corresponding feature cell in the bounding box $\mathbf{\tilde{b}}$ that lies within an ellipse centered in it. The impact of this spatial prior and the underlying assumption is further analyzed through ablation experiments in LABEL:ssec:ablation.
Figure 1: Overview of CountingDINO. Given an image $I$ and $N$ exemplar boxes, we extract features using the DINO-based visual backbone and apply each exemplar as a convolutional kernel over the image feature map to obtain similarity maps. These are aggregated, normalized into a density map using spatial priors, and thresholded before integration to produce the final count.
3.2 Similarity Map Generation
In this stage, we create a 2D similarity map where each value reflects how closely a local image region matches the exemplar in appearance and structure. To obtain this, we use each pooled exemplar feature $\mathbf{R}^{\mathbf{\tilde{b}}}$ as a convolutional kernel on the map $\mathbf{V}$, i.e., $\mathbf{S}^{\mathbf{\tilde{b}}}=\text{Conv2D}(\mathbf{V},\mathbf{R}^{\mathbf{\tilde{b}}})\in\mathbb{R}^{\bar{L}\times\bar{V}}$, where $\bar{L}$ and $\bar{V}$ are the spatial dimensions of the output map.
We repeat this process for each of the $N$ exemplar bounding boxes $\{\mathbf{\tilde{b}}_{i}\}_{i=1}^{N}$, resulting in $N$ similarity maps ${\mathbf{S}_{1},\mathbf{S}_{2},\dots,\mathbf{S}_{N}}$. Finally, we average the aligned maps to obtain a single aggregated similarity map $\mathbf{S}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{S}_{i}$, which reflects the overall likelihood of object occurrence across all exemplars.
3.3 Density Map Normalization
The aggregated similarity map S well highlights regions likely to contain object instances. However, we seek a similarity response that integrates to 1 over each object instance, effectively transforming the similarity map into a density map from which the total count can be estimated via integration.
In exemplar-based counting, we can leverage the $N$ exemplar bounding boxes provided as input to properly estimate a normalization factor. In fact, we know that the area occupied by the $N$ exemplars should count up to exactly $N$. To this aim, we reuse the elliptical masks $\{\mathbf{M}^{\mathbf{\tilde{b}}_{i}}\}_{i=1}^{N}$ introduced in section 3.1 and compute the normalization factor $z$ as follows:
$z=\frac{1}{N}\sum_{(x,y)\in(\bar{L}\times\bar{V})}\mathbf{M}_{xy}\odot\text{minmax}(\mathbf{S})_{xy},$ (1)
where the global weighting mask $\mathbf{M}$ is obtained by properly rescaling the masks $\{\mathbf{M}^{\mathbf{\tilde{b}}_{i}}\}_{i=1}^{N}$ and accumulating them into a $\bar{L}\times\bar{V}$ map initialized with all zeros. Notice that, to ensure the normalization relies only on positive values, we first apply minmax normalization to $\mathbf{S}$, rescaling it to the range $[0,1]$.
The final normalized map is obtained by dividing the minmax-ed activation map by the computed scalar factor $z$:
$\hat{\mathbf{S}}=\frac{\text{minmax}(\mathbf{S})}{z}$ (2)
This ensures the total response across the $N$ exemplar regions sums to $N$, making the response per object approximate a unit mass. Consequently, the normalized map $\hat{\mathbf{S}}$ serves as a density estimate, with the total count obtained by integrating it over the spatial dimensions.
3.4 Thresholding density map contributions
The similarity map highlights object regions but also includes low activations in the background, which can cause overcounting during integration. To address this, we apply a thresholding step to suppress low-activation areas.
In order to find a reasonable threshold, we notice that the normalization criteria detailed in Eqs. 1 and 2 induce a mean per-patch unit count, estimated on the set $\mathcal{B}$ of exemplars, of $u(\mathcal{B})=\frac{|\mathcal{B}|}{\sum_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})}$ – where $\text{area}(\cdot)$ is a function returning the area, expressed in the number of patches, for the given bounding box. The resulting term $u(\mathcal{B})$ is the mean unit count value that a patch carries if it is part of a relevant object. This means that it is likely that patches carrying less value than $u(\mathcal{B})$ pertain to a non-relevant object and may be filtered out.
However, notice that $u(\mathcal{B})$ assumes its minimum value when fed with the set =$\mathcal{B^{*}}=\{\operatorname*{arg\,max}_{\mathbf{b}\in\mathcal{B}}\text{area}(\mathbf{b})\}$ – i.e. when computed on the set composed solely of the largest exemplar bounding box. In cases different from this, we are also inevitably filtering out contributions from the other exemplars bounding boxes that we know are valid. For this reason, we set the threshold to be exactly the unit count computed on the largest region, i.e., $\tau=u(\mathcal{B}^{*})$.
This process ensures that only the relevant activations are considered, improving the accuracy of the object count estimate.
3.5 Increasing Spatial Resolution
The above-presented method generates consistent results when the spatial resolution of DINO is enough to capture no more than an object per visual patch, which happens when the objects are not too small. Therefore, instead of processing the entire image at once, we divide the input image into four non-overlapping and evenly-shaped quadrants and independently process each sub-image with the backbone $\psi_{\text{v}}$. In general, we divide the image into $4^{k}$ different quadrants, where $k\in\mathbb{N}$ defines the resolution level – i.e., how many times we recursively partition each quadrant in its four sub-quadrants.
After extracting features from each quadrant, we spatially reassemble the four feature maps into a single unified feature tensor by stitching them according to their original spatial layout. This yields a higher-resolution feature map $\mathbf{V}\in\mathbb{R}^{2^{k}L\times 2^{k}V\times D}$, effectively increasing the granularity of the representation and enabling more precise localization of small objects.
This modification integrates seamlessly into the pipeline, as all subsequent steps – ROI-Align, similarity computation, and normalization – remain unchanged. We found $k=2$ to work well in the presented domain, although we experiment with other values in LABEL:ssec:ablation.
4 Experimental Evaluation
4.1 Datasets and Metrics
We evaluate our method on the FSC-147 dataset [DBLP:conf/cvpr/RanjanSNH21], which consists of 6,135 images belonging to 147 object classes, each accompanied by three exemplar bounding boxes. We evaluate counting performance using two standard metrics: Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) [DBLP:conf/eccv/ArtetaLZ16, DBLP:conf/nips/LempitskyZ10, DBLP:conf/cvpr/RanjanSNH21], defined as $MAE=\frac{1}{n}\sum_{i=1}^{n}|c_{i}-\hat{c}_{i}|$ and $RMSE=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(c_{i}-\hat{c}_{i})^{2}}$, where $c_{i}$ and $\hat{c}_{i}$ are the ground truth and predicted counts for the $i$-th image.
4.2 Baseline
Since CountingDINO is the first training-free methodology for exemplar-based CAC using an unsupervised backbone, we design a comparison baseline by adapting CutLER [wang2023cut] – an unsupervised object detector that leverages pseudo-labels derived from DINO features. As CutLER is not designed for exemplar-based matching, we adapt it to our setting as follows. We first apply CutLER to the target image to obtain a set of detected object bounding boxes. To filter these detections and retain only those corresponding to the exemplar object category, we extract visual features using DINO – specifically, DINO ViT-B/8, the same backbone used during CutLER’s training. For each exemplar bounding box, we apply ROI pooling to obtain its feature representation, and average these representations to form a prototype feature vector representing the exemplar class. We then extract features from the detected bounding boxes and compute their similarity to the prototype vector. Detections with a similarity score above a threshold (set to 0.5) are retained, and their count provides the final prediction.